Solve for 3 3 - 3 x + 2 X – 110 10 4 Solve the rational equation: + - = m 7 m Answer: m = Preview Enter your answer as an integer or rational number in th14 Solve the rational equation: a – + 5 a Answer: a = Enter your answers as a list of integers or rational numbers in the for
Answer #1

SOLUTION :


1.


3/(x - 1) = - 3/(x+2)


Multiply by (x - 1)(x + 2) to both sides :

=> 3(x + 2)  = - 3(x - 1)

=> 3x + 6 = - 3x + 3


Add 3x to, and subtract 6 from both sides :

=> 6x = - 3 

Divide both sides by 6 :

=> x = - 3/6

=> x = - 1/2 (ANSWER).


2.


- 10/m + 10/7 = 4/m 


Multiply by 7m , both sides :

=> - 70 + 10m = 28

Add 70 to booth sides :

=> 10m = 28 + 70 = 98

=> m = 98/10 

=> m = 9.8 (ANSWER).

answered by: Tulsiram Garg